OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (2, -2) with 0° ≤ θ < 360°.

OpenStudy (agent0smith):

Polar coordinates means you need r (the magnitude) and the angle θ. Your point (x, y) is (2, -2) r^2 = x^2 + y^2 Find θ using tan θ = y/x (then use inverse tan to find θ)

OpenStudy (agent0smith):



OpenStudy (anonymous):

i got r=0 and θ 45

OpenStudy (agent0smith):

r^2 = x^2 + y^2... i think you may not have squared them, since (-2)^2 = 4 45 is correct.

OpenStudy (agent0smith):

-45 is the angle, not +45.

OpenStudy (anonymous):

i still get 0

OpenStudy (agent0smith):

r^2 = 2^2 + (-2)^2 =

OpenStudy (anonymous):

8 but if i square it 4?

OpenStudy (agent0smith):

No it's just square root of 8, but you can simplify \[\large r = \sqrt 8 = \sqrt 4 \sqrt 2 = 2 \sqrt 2\]

OpenStudy (anonymous):

(2 square root of 2, 45°), (-2 square root of 2, 225°)

OpenStudy (agent0smith):

The angle is -45, so we can add 360 to that to get 315 so one is (2 square root of 2, 315°)

OpenStudy (agent0smith):

the other, we can make r = -2sqrt2, and subtract 180 from the angle 315

OpenStudy (anonymous):

wait so am i right

OpenStudy (agent0smith):

(2 square root of 2, 45°), (-2 square root of 2, 225°) aren't correct

OpenStudy (agent0smith):

well the second is, first isn't

OpenStudy (anonymous):

(2 square root of 2, 135°), (-2 square root of 2, 315°)

OpenStudy (agent0smith):

Wait, no, we had one already: (2 square root of 2, 315°)

OpenStudy (agent0smith):

all you have to do is make r negative, and subtract 180 from the angle, that's your other

OpenStudy (anonymous):

(2 square root of 2, 315°), (-2 square root of 2, 135°)

OpenStudy (agent0smith):

Good :)